Steve is 25 years older than his son John. In ten years, Steve will be twice John's age. How old are Steve and John now? Write a system of equations to model the situation.
Answer by stanbon(75887)

Steve is 25 years older than his son John. In ten years, Steve will be twice John's age. How old are Steve and John now? Write a system of equations to model the situation.
------
Equation:
s = j + 25
(s+10) = 2(j+10)
-------
Substitute for "s" and solve for "j":
(j+25+10) = 2j+20
---
j+35 = 2j+20
j = 15 (John's age now)
------
Solve for "s":
s = j+25
s = 15+25 = 40 (Steve's age now)
===============
Cheers,
Stan H.